# Revisiting the Predictability of Performative, Social Events

### v1

@@ 1 Introduction
In this work, we revisit this question using modern mathematical tools. Framing our analysis in the language of performative prediction – a recent learning-theoretic framework introduced by Perdomo et al. [2020] that formalizes the causal aspects of social prediction – we establish the following result. One can always efficiently find predictions that actively shape outcomes and simultaneously satisfy rigorous validity guarantees like multi-calibration [Hébert-Johnson et al., 2018] or outcome indistinguishability [Dwork et al., 2021]. Moreover, the statistical and computational complexity of finding these predictors is, in many cases, just as easy as in supervised learning, where the distribution is static.
However, while these rigorously calibrated predictions are always achievable, they are not always desirable. By only focusing on the degree to which predictions match outcomes, these historical calibration desiderata ignore that one can achieve better performance by actively steering social patterns and shaping the underlying distribution. Through simple examples, we show that these prediction rules can lead to poor equilibria where predictions exactly match the conditional distribution over outcomes – and hence are perfectly multicalibrated – while simultaneously explaining none of the variance. We conclude by discussing avenues forward.
@@ 1.1 Overview of Technical Results
We cast this problem in the language of performative prediction [Perdomo et al., 2020], a recent learning theoretic framework that formalizes how data in social contexts is not static, but rather a function of the published predictor $f$. The key conceptual device of the framework is the distribution map ${\cal D}(\cdot)$, a function which maps predictors $f$ to distributions over feature, outcome pairs $(x,y)$. It captures how deploying different predictors induces different distributions.
@@
Here, ${\cal C}$ parametrizes the degree to which predictions match outcomes. If ${\cal C}$ consists of just the constant 1 function, we recover the guarantee considered in Simon [1954] and Grunberg and Modigliani [1954] that $\operatorname*{\mathbb{E}}_{{\cal D}(f)}[y]=\operatorname*{\mathbb{E}}_{{\cal D}(f)}[f(x)]$ overall. If ${\cal C}$ consists of all bounded, measurable functions, then $f(x)$ must be equal to the conditional distribution over outcomes. We can, of course, interpolate between these two extremes by varying ${\cal C}$.
While the guarantee in Definition 1.1 is known to be achievable in supervised learning where ${\cal D}(f)={\cal D}(f^{\prime})$ for all $f$ and $f^{\prime}$, the situation is substantially more complicated in this performative setting. Predictions can, in general, be self-negating and outcomes can move away from the published prediction $p\sim f(x)$. Given this possibility, it is not immediately obvious that performatively calibrated $f$ even exist (that is $f$ such that $\operatorname*{\mathbb{E}}_{{\cal D}(f)}[f(x)]=\operatorname*{\mathbb{E}}_{{\cal D}(f)}[y]$). The situation is further complicated by the fact that the distribution map ${\cal D}(\cdot)$ is unknown to the learner. They can only deploy a predictor $f$ and observe the induced samples $(x,p,y)\sim{\cal D}(f)$.
Our first result shows that this guarantee Definition 1.1 is, in fact, efficiently achievable both statistically and computationally, thereby providing a modern learning-theoretic answer to the question posed by Morgenstern, Simon, and others.
@@
$\big{|}\sum_{i=t}^{T}c(x_{t},p_{t})(y_{t}-p_{t})\big{|}\leqslant\mathsf{Regret}(T)\text{ for all }c\in{\cal C},$ (1)
@@
The result establishes an efficient reduction from our main problem to an online multicalibration problem for which numerous algorithms now exist [Vovk, 2007, Foster and Kakade, 2006, Dwork et al., 2024, Gupta et al., 2022, Okoroafor et al., 2025]. In particular, these algorithms can efficiently achieve the online guarantee in Equation 1 for various rich classes of functions ${\cal C}$ such as low-degree polynomials, decision trees, or, more generally, any ${\cal C}$ that is (weak agnostically) learnable or that belongs to a reproducing kernel Hilbert space. Moreover, these algorithms have $\sqrt{T}$ regret, implying that $f$ achieves the optimal ${\cal O}(n^{-1/2})$ rate for this problem. The reduction is conceptually simple, yields tight bounds on the generalization error, and—unlike the bulk of work in performative prediction —requires no Lipschitzness conditions on the distribution map ${\cal D}(\cdot)$. That is, the conditional distribution over outcomes ${\cal D}_{y}(x,p)$ can be any discontinuous, non-smooth function of the forecast $p$. We only need to assume that outcomes lie in a bounded range.
@@
However, leaving the choice of loss function aside and restricting ourselves to predictive accuracy as the main criterion, prediction rules satisfying the historical desiderata formalized in Definition 1.1 can be arbitrarily poor. We show that it is possible for a predictor $f$ to be performatively multicalibrated with respect to all (continous) functions $c(x,p)$ while simultaneously explaining none of the variance in outcomes. We state the following result in terms of the core solution concepts from the performative prediction framework —performative stability and performative optimality — whose definitions we review below. [^2] The solution concepts of performative stability and optimality are defined with respect to general loss functions $\ell(x,y;h)$ that can capture a wide variety of higher level objectives. For instance, we can choose to have losses which encourage forecasts to match outcomes, $\ell(x,y;h)=(h(x)-y)^{2}$, or losses which encourage the induced distributions ${\cal D}(f)$ towards a particular outcome $\ell(x,y;h)=-y$. See Kim and Perdomo [2023]. Since the goal here is to understand the relationship between from Definition 1.1 and predictive accuracy in performative contexts, we specialize the definitions to the case of squared loss for simplicity.
@@ 1.2 Related Work
Our results build on the recent lines of work on multicalibration [Hébert-Johnson et al., 2018] and outcome indistinguishability [Dwork et al., 2021], further extending the foundations of these ideas to non-supervised learning settings. Our proofs heavily rely on new algorithmic insights from online calibration, pioneered by Foster and Vohra [1998], and extended by [Foster and Kakade, 2006, Kakade and Foster, 2008, Vovk, 2007, Gupta et al., 2022] and others. Within this broad literature, our work is most closely related to Kim and Perdomo [2023], who used these tools developed in Gopalan et al. [2023] to compute performatively optimal (not stable) predictors for a restricted version of the outcome performativity setting. In their setup, outcomes are binary and their conditional distribution is a function of $x$ and a discrete decision $\hat{y}$ belonging to a finite set. Our setup generalizes theirs since both outcomes $y$ and forecasts $p$ can be real-valued. [^4] In particular, we can recover theirs by for instance requiring $p\in[0,1]$ and partitioning the unit interval into sets $P_{i}$ such that if $p\in P_{i}$ then the learner chooses action $\hat{y}_{i}$.
@@ 2 Preliminaries
Throughout our work, we assume that data $(x,p,y)\sim{\cal D}(f)$ is generated according to the following process. The learner publishes a predictor $f:\mathcal{X}\rightarrow\Delta([0,1])$ mapping features $x$ to a distribution $f(x)$ over the unit interval $[0,1]$. Features $x$ are drawn i.i.d from a fixed distribution ${\cal D}_{x}$ over an arbitrary set $\mathcal{X}$, predictions $p\in[0,1]$ are sampled from $f$, $p\sim f(x)$. Outcomes $y$ are sampled from ${\cal D}_{y}(x,p)$ where ${\cal D}_{y}(x,p)$ is any distribution over the unit interval $[0,1]$. Our results apply to any ${\cal D}_{y}(x,p)$ supported on a bounded interval by simply rescaling.
@@
Calibration is the sine qua non definition of validity for a probabilistic forecast [Dawid, 1985]. For binary $y$, a predictor is calibrated if conditional on $f(x)=v$ for $v\in[0,1]$, the outcome $y$ occurs a $v$ fraction of the time, $\mathrm{Pr}[y=1\mid f(x)=v]=v$. Multicalibration [Hébert-Johnson et al., 2018] is a strengthening of calibration, requiring that $f(x)=y$, not just overall, but also once we condition on $x$ belonging to any set $G$ in a collection ${\cal G}$: $\mathrm{Pr}[y=1\mid f(x)=v,x\in G]=v$ for all $G\in{\cal G}$. This condition is equivalent, up to a normalization factor of $\mathrm{Pr}[f(x)=v,x\in G]$, to the one we wrote in Definition 1.1 since letting $c_{G,v}(x,p)=1\{x\in G,p=v\}$, we can write:
@@
Conversely, a predictor $f$ is outcome indistinguishable (or OI) [Dwork et al., 2021] if it establishes a generative model that cannot be falsified on the basis of a prespecified collection of tests or distinguishers $A\in{\cal F}\subseteq\{\mathcal{X}\times[0,1]\times[0,1]\rightarrow\mathbb{R}\}$. Simplifying our discussion to the case of a binary outcome, each distinguisher takes features $x$, predictions $p$, and an outcome $y$ and outputs 1 or 0 (i.e. is this a real outcome/prediction for $x$). A predictor is OI if all the distinguishers in the collection ${\cal F}$ behave the same when given the true outcome $y\sim{\cal D}(f)$ versus an outcome sampled from the model $\tilde{y}\sim\mathrm{Ber}(p),p\sim f(x)$. Extended to the performative context, $f$ is outcome indistinguishable with respect to ${\cal F}$ if for all $A\in{\cal F}$
$\operatorname*{\mathbb{E}}_{(x,p,y)\sim{\cal D}(f)}A(x,p,y)=\operatorname*{\mathbb{E}}_{x\sim{\cal D}_{x}p\sim f(x),\tilde{y}\sim\mathrm{Ber}(p)}A(x,p,\tilde{y})$ (4)
@@ 3 Algorithmic Results
Recall that the goal is to find a prediction rule $f$ satisfying the following indistinguishability guarantee from Definition 1.1 such that for all $c\in{\cal C}$,
@@
for all $c\in{\cal C}$ where $\mathsf{Regret}_{{\cal A}}(T):\mathbb{N}\rightarrow\mathbb{R}_{\geqslant 0}$, is $o(T)$.
Algorithms that achieve this guarantee date back to the work Vovk [2007], Vovk et al. [2005a] and Foster and Kakade [2006] (they refer to it under different names like resolution or just calibration). Following the work of Hébert-Johnson et al. [2018] introducing multicalibration, there has been a flurry of recent papers introducing new algorithms for this problem [Dwork et al., 2024, Gupta et al., 2022, Garg et al., 2024, Okoroafor et al., 2025]. We will make use of these procedures when instantiating our general results. As a final preliminary step, we define the online-to-batch procedure we use in our analysis:
@@
This style of online-to-batch conversion, where one uniformly randomizes over all previous predictors, is standard in the online learning literature and is the typical starting point when converting online algorithms to batch learners (see, e.g., Gupta et al. [2022], Okoroafor et al. [2025]). The main difference in this construction is that samples $(x_{t},y_{t})$ are not drawn i.i.d from a fixed distribution ${\cal D}$, but rather from the distribution $(x_{t},y_{t})\sim{\cal D}(f_{t})$ induced by the predictions. Note that while each of the individual functions $f_{t}\in\{f_{1},\dots,f_{t}\}$ in the transcript are deterministic, the batch predictor is randomized since it first mixes over the choice of $f_{t}$. With these definitions out of the way, we can now state the main theorem for this section:
@@
    Any Finite Collection. Let ${\cal C}\subseteq\{\mathcal{X}\times[0,1]\rightarrow[-1,1]\}$ be any finite set of functions $c(x,p)$ that are continuous in the forecast $p$. Then, there exists a choice of kernel, such that the K29 algorithm is online multicalibrated with respect to ${\cal C}$ at a rate bounded by $\sqrt{T|{\cal C}|}$. Consequently, the batch version $f$ trained on $n$ rounds of interaction satisfies:
@@
    (Low-Degree Boolean Functions). Define $\mathcal{X}=\{0,1\}^{d}$, and ${\cal C}_{\mathrm{LowDeg}}\subseteq\{\mathcal{X}\rightarrow[-1,1]\}$ to be the set of Boolean functions of degree $s$. That is, those that can be written as,
@@
The proofs of all these statements follow exactly the same simple template. We simply restate known regret bounds for the various algorithms that appear in the literature and then plug these into Theorem 3.4. ∎
The corollary illustrates how one can efficiently find prediction rules $f$ that are performatively multicalibrated with respect to common classes of functions, for instance, linear functions, low-degree polynomials, or even random forests. [^7] In more detail, the claim about low-degree polynomials and random forests follows from part $b$ of the corollary. It is well known that degree $s$ polynomials and depth $s$ regression trees are low-degree boolean functions. We refer the reader to [O’Donnell, 2014] for further background. Furthermore, a random forest $h(x)$ is simply a sum of regression trees $c_{j}$, $h(x)=\sum_{j}\alpha_{j}c_{j}(x)$. Note that the online guarantee is closed with respect to addition. If $\sum_{t=1}^{T}c_{j}(x_{t})(y_{t}-p_{t})\leqslant o(T)$, then $|\sum_{t=1}^{T}\sum_{j}\alpha_{j}c_{j}(x_{t})(y_{t}-p_{t})|\leqslant\sum_{j}|\alpha_{j}\sum_{t=1}^{T}c_{j}(x_{t})(y_{t}-p_{t})|\leqslant\sum_{j}|\alpha_{j}||\sum_{t=1}^{T}c_{j}(x_{t})(y_{t}-p_{t})|$ which is $o(T)$ if $\sum_{j}|\alpha_{j}|\leqslant{\cal O}(1)$. One can even efficiently guarantee indistinguishability with respect to any polynomially sized collection ${\cal C}$ (as long as the functions $c\in{\cal C}$ are each efficiently computable). This, in particular, implies performative indistinguishability with respect to rich classes of functions, like deep neural networks. These hold without making any Lipschitzness assumptions on the distribution map ${\cal D}(\cdot)$. Furthermore, the runtime of the procedures is, up to small polynomial factors, no different than that of algorithms achieving solving the analogous guarantee in supervised learning settings.

### v2

@@ 1 Introduction
In this work, we revisit this question using modern mathematical tools. Framing our analysis in the language of performative prediction – a recent learning-theoretic framework introduced by Perdomo et al. [2020] that formalizes the causal aspects of social prediction – we establish the following result. One can always efficiently find predictions that actively shape the data distribution over outcomes and simultaneously satisfy rigorous validity guarantees like multi-calibration [Hébert-Johnson et al., 2018] or outcome indistinguishability [Dwork et al., 2021]. Moreover, the statistical and computational complexity of finding these predictors is, in many cases, just as easy as in supervised learning, where the distribution is static.
However, while these rigorously calibrated predictions are always achievable, they are not always desirable. The fact that predictions are performative and influence the likelihood of future events invalidates naive extensions of traditional solution concepts inherited from supervised learning like calibration. Through simple examples, we show that these prediction rules can lead to poor equilibria where predictions exactly match the conditional distribution over outcomes – and hence are perfectly multicalibrated – while simultaneously explaining none of the variance in outcomes. We conclude by discussing avenues forward.
@@ 1.1 Overview of Technical Results
We cast this problem in the language of performative prediction [Perdomo et al., 2020], a recent learning theoretic framework that formalizes how data in social contexts is not static, but rather a function of the published predictor $f$. The key conceptual device of the framework is the notion of the distribution map ${\cal D}(\cdot)$, a function which maps predictors $f$ to distributions over feature, outcome pairs $(x,y)$. It captures how different predictors induces different distributions.
@@
Here, ${\cal C}$ parametrizes the degree to which predictions match outcomes. If ${\cal C}$ consists of just the constant 1 function, we recover the guarantee considered in Simon [1954] and Grunberg and Modigliani [1954] that $\operatorname*{\mathbb{E}}_{{\cal D}(f)}[y]=\operatorname*{\mathbb{E}}_{{\cal D}(f)}[f(x)]$ overall. If ${\cal C}$ consists of all bounded, measurable functions, then $f(x)$ must be equal to the conditional distribution over outcomes, $f(x)=\operatorname*{\mathbb{E}}_{{\cal D}(f)}[y|x]$ for all $x$. We can interpolate between these two extremes by varying ${\cal C}$.
While the guarantee in Definition 1.1 is known to be achievable in supervised learning where ${\cal D}(f)={\cal D}(f^{\prime})$ for all $f$ and $f^{\prime}$, the situation is substantially more complicated in this performative setting. Predictions can, in general, be self-negating and outcomes can move away from the published prediction $p\sim f(x)$. Given this possibility, it is not immediately obvious that performatively calibrated $f$ even exist (that is a predictor $f$ such that $\operatorname*{\mathbb{E}}_{{\cal D}(f)}[f(x)]=\operatorname*{\mathbb{E}}_{{\cal D}(f)}[y]$) without making strong regularity assumptions on ${\cal D}(\cdot)$. The situation is further complicated by the fact that the distribution map ${\cal D}(\cdot)$ is unknown to the learner. They can only deploy a predictor $f$ and observe the induced samples $(x,p,y)\sim{\cal D}(f)$.
Our first result shows that this guarantee Definition 1.1 is, in fact, efficiently achievable both statistically and computationally, thereby providing a modern learning-theoretic answer to the question posed by Morgenstern, Simon, and others. Unlike the bulk of work in performative prediction [Perdomo et al., 2020, Mendler-Dünner et al., 2020], this result requires no smoothness or continuity assumptions on the outcome performative distribution map ${\cal D}(\dot{)}$ other than the condition that outcomes $y$ lie in a bounded range.
@@
$\big{|}\sum_{i=t}^{T}c(x_{t},p_{t})(y_{t}-p_{t})\big{|}\leqslant\mathsf{Regret}(T)\text{ for all }c\in{\cal C}.$ (1)
@@
The result establishes an efficient reduction from our main problem to an online multicalibration problem for which numerous algorithms now exist [Vovk, 2007, Foster and Kakade, 2006, Dwork et al., 2025, Gupta et al., 2022, Okoroafor et al., 2025, Perdomo and Recht, 2025]. In particular, these algorithms can efficiently achieve the online guarantee in Equation 1 for various rich classes of functions ${\cal C}$ such as low-degree polynomials, decision trees, or, more generally, any ${\cal C}$ that is (weak agnostically) learnable or that belongs to a reproducing kernel Hilbert space. Moreover, these algorithms have $\sqrt{T}$ regret, implying that $f$ achieves the optimal ${\cal O}(n^{-1/2})$ rate for this problem. The reduction is conceptually simple, yields tight bounds on the generalization error, and requires no Lipschitzness conditions on the distribution map ${\cal D}(\cdot)$. That is, the conditional distribution over outcomes ${\cal D}_{y}(x,p)$ can be any discontinuous, non-smooth function of the forecast $p$. We only need to assume that outcomes lie in a bounded range.
@@
However, leaving the choice of loss function aside and restricting ourselves to predictive accuracy as the main criterion, prediction rules satisfying the historical desiderata formalized in Definition 1.1 can be arbitrarily poor. We show that it is possible for a predictor $f$ to be performatively multicalibrated with respect to all (continuous) functions $c(x,p)$ while simultaneously explaining none of the variance in outcomes. We state the following result in terms of the core solution concepts from the performative prediction framework —performative stability and performative optimality — whose definitions we review below. [^2] The solution concepts of performative stability and optimality are defined with respect to general loss functions $\ell(x,y;h)$ that can capture a wide variety of higher level objectives. For instance, we can choose to have losses which encourage forecasts to match outcomes, $\ell(x,y;h)=(h(x)-y)^{2}$, or losses which encourage the induced distributions ${\cal D}(f)$ towards a particular outcome $\ell(x,y;h)=-y$. See Kim and Perdomo [2023]. Since the goal here is to understand the relationship between Definition 1.1 and predictive accuracy in performative contexts, we specialize the definitions to the case of squared loss for simplicity.
@@
Note that our intuition from supervised learning is exactly reversed. In supervised learning, where $(x,y)$ are drawn from a fixed distribution ${\cal D}$, if $f(x)=\operatorname*{\mathbb{E}}[y\mid x]$ for all $x$, then $\operatorname*{\mathbb{E}}_{{\cal D}}(y-f(x))^{2}\leqslant\operatorname*{\mathbb{E}}_{{\cal D}}(y-h(x))^{2}$ for any function $h$. In performative prediction, where $f$ influences the distribution over $(x,y)$, the result shows that it is possible for $f(x)=\operatorname*{\mathbb{E}}_{{\cal D}(f)}[y\mid x]$ for every $x$, yet $\operatorname*{\mathbb{E}}_{{\cal D}(f)}(y-f(x))^{2}\geqslant\operatorname*{\mathbb{E}}_{{\cal D}(h)}(y-h(x))^{2}$. The inequality has now been flipped because of performativity.
@@ 1.2 Related Work
Our results also build on the recent lines of work on multicalibration [Hébert-Johnson et al., 2018] and outcome indistinguishability [Dwork et al., 2021], further extending the foundations of these ideas to non-supervised learning settings. Our proofs heavily rely on new algorithmic insights from online calibration, pioneered by Foster and Vohra [1998], and extended by [Foster and Kakade, 2006, Kakade and Foster, 2008, Vovk, 2007, Gupta et al., 2022] and others. Within this broad literature, our work is most closely related to Kim and Perdomo [2023], who used these tools developed in Gopalan et al. [2023] to compute performatively optimal (not stable) predictors for a restricted version of the outcome performativity setting. In their setup, outcomes are binary and their conditional distribution is a function of $x$ and a discrete decision $\hat{y}$ belonging to a finite set. Our setup generalizes theirs since both outcomes $y$ and forecasts $p$ can be real-valued.
@@ 2 Preliminaries
Throughout our work, we assume that data $(x,p,y)\sim{\cal D}(f)$ is generated according to the following process. The learner publishes a predictor $f:\mathcal{X}\rightarrow\Delta([0,1])$ mapping features $x$ to a distribution $f(x)$ over the unit interval $[0,1]$. Features $x$ are drawn i.i.d from a fixed distribution ${\cal D}_{x}$ over an arbitrary set $\mathcal{X}$, predictions $p\in[0,1]$ are sampled from $f(x)$, $p\sim f(x)$. Outcomes $y$ are sampled from ${\cal D}_{y}(x,p)$ where ${\cal D}_{y}(x,p)$ is any distribution over the unit interval $[0,1]$. Our results apply to any ${\cal D}_{y}(x,p)$ supported on a bounded interval by rescaling.
@@
Calibration is the sine qua non definition of validity for a probabilistic forecast [Dawid, 1985]. For binary $y$, a predictor is calibrated if conditional on $f(x)=v$ for $v\in[0,1]$, the outcome $y$ occurs a $v$ fraction of the time, $\mathrm{Pr}[y=1\mid f(x)=v]=v$.
Multicalibration [Hébert-Johnson et al., 2018] is a strengthening of calibration, requiring that $f(x)=y$, not just overall, but also once we condition on $x$ belonging to any set $G$ in a collection ${\cal G}$: $\mathrm{Pr}[y=1\mid f(x)=v,x\in G]=v$ for all $G\in{\cal G}$. This condition is equivalent, up to a normalization factor of $\mathrm{Pr}[f(x)=v,x\in G]$, to the one we wrote in Definition 1.1 since letting $c_{G,v}(x,p)=1\{x\in G,p=v\}$, we can write:
@@
Conversely, a predictor $f$ is outcome indistinguishable (or OI) [Dwork et al., 2021] if it establishes a generative model that cannot be falsified on the basis of a prespecified collection of tests or distinguishers $A\in{\cal F}\subseteq\{\mathcal{X}\times[0,1]\times[0,1]\rightarrow\mathbb{R}\}$. Simplifying our discussion to the case of a binary outcome, each distinguisher takes features $x$, predictions $p$, an outcome $y$ and outputs 1 or 0 (i.e. is this a real outcome/prediction for $x$). A predictor is OI with respect to ${\cal F}$ if all the distinguishers in the collection ${\cal F}$ behave the same when given the true outcome $y\sim{\cal D}(f)$ versus an outcome sampled from the model $\tilde{y}\sim\mathrm{Ber}(p),p\sim f(x)$. Extended to the performative context, $f$ is outcome indistinguishable with respect to ${\cal F}$ if for all $A\in{\cal F}$
$\operatorname*{\mathbb{E}}_{(x,p,y)\sim{\cal D}(f)}A(x,p,y)\approx\operatorname*{\mathbb{E}}_{x\sim{\cal D}_{x}p\sim f(x),\tilde{y}\sim\mathrm{Ber}(p)}A(x,p,\tilde{y})$ (4)
@@ 3 Algorithmic Results
Recall that the goal is to find a prediction rule $f$ satisfying the following indistinguishability guarantee from Definition 1.1. For all $c\in{\cal C}$,
@@
for all $c\in{\cal C}$ where $\mathsf{Regret}_{{\cal A}}(T):\mathbb{N}\rightarrow\mathbb{R}_{\geqslant 0}$ is $o(T)$.
Algorithms that achieve this guarantee date back to the work Vovk [2007], Vovk et al. [2005a] and Foster and Kakade [2006] (they refer to it under different names like resolution or just calibration). Following the work of Hébert-Johnson et al. [2018] introducing multicalibration, there has been a flurry of recent papers introducing new algorithms for this problem [Dwork et al., 2025, Gupta et al., 2022, Garg et al., 2024, Okoroafor et al., 2025]. We will make use of these procedures when instantiating our general results. As a final preliminary step, we define the online-to-batch procedure we use in our analysis:
@@
This style of online-to-batch conversion where one uniformly randomizes over all previous predictors is standard in the online learning literature and is the typical starting point when converting online algorithms to batch learners (see, e.g., Gupta et al. [2022], Okoroafor et al. [2025]). The main difference in this construction is that samples $(x_{t},y_{t})$ are not drawn i.i.d from a fixed distribution ${\cal D}$, but rather from the distribution $(x_{t},y_{t})\sim{\cal D}(f_{t})$ induced by the predictions. Note that while each of the individual functions $f_{t}\in\{f_{1},\dots,f_{t}\}$ in the transcript are deterministic, the batch predictor is randomized since it first mixes over the choice of $f_{t}$. With these definitions out of the way, we can now state the main theorem for this section:
@@
    Any Finite Collection. Let ${\cal C}\subseteq\{\mathcal{X}\times[0,1]\rightarrow[-1,1]\}$ be any finite set of functions $c(x,p)$ that are continuous in the forecast $p$. Then, there exists an explicit choice of kernel, such that the K29 algorithm is online multicalibrated with respect to ${\cal C}$ at a rate bounded by $\sqrt{T|{\cal C}|}$. Consequently, the batch version $f$ trained on $n$ rounds of interaction satisfies:
@@
    Low-Degree Boolean Functions. Define $\mathcal{X}=\{0,1\}^{d}$, and ${\cal C}_{\mathrm{LowDeg}}\subseteq\{\mathcal{X}\rightarrow[-1,1]\}$ to be the set of Boolean functions of degree $s$. That is, those that can be written as,
@@
The proofs of all these statements follow exactly the same simple template. We simply restate known regret bounds for the various algorithms that appear in the literature and then plug these into Theorem 3.4. While we omit the precise form of each kernel from the theorem statement, the interested reader can find them in the appendix. ∎
The corollary illustrates how one can efficiently find prediction rules $f$ that are performatively multicalibrated with respect to common classes of functions, for instance, linear functions or low-degree polynomials. One can even efficiently guarantee indistinguishability with respect to any polynomially-sized collection ${\cal C}$ (as long as the functions $c\in{\cal C}$ are each efficiently computable). This, in particular, implies performative indistinguishability with respect to rich classes of functions, like deep neural networks. These hold without making any Lipschitzness assumptions on the distribution map ${\cal D}(\cdot)$. Furthermore, the runtime of the procedures is, up to small polynomial factors, no different than that of algorithms achieving solving the analogous guarantee in supervised learning settings.